# Progressive Contrast Registration for High-Fidelity Bidirectional Photoacoustic Microscopy Alignment

### v1

@@ 3.2 Comparison with State-of-the-Art
fig. 2 presents qualitative comparison on representative test samples.
Figure 2: Qualitative comparison on a representative 26K test sample. Top row: merged full frame (odd + registered even interleaved) in hot colormap; smooth appearance indicates good alignment. Bottom row: magenta/green overlay where odd columns are magenta and registered even columns are green; well-aligned regions appear gray, misaligned regions show colored fringes. Traditional methods (SIFT, Demons, OF, SyN) and deformation-based methods (VM, TM) exhibit visible artifacts, while our method achieves near-perfect alignment with no visible fringes.
3.3 Temporal Consistency Evaluation
Beyond single-frame alignment quality, we evaluate temporal consistency across consecutive frames—a critical requirement for longitudinal OR-PAM studies. Using the TNCC and TNCG metrics defined in section 3.1, we assess whether each registration method preserves the natural temporal dynamics when registered even columns are merged with odd columns to form full frames.
Table 2: Temporal consistency evaluation on OR-PAM-Reg-Temporal-26K (2,691 consecutive frame pairs). NCC and SSIM measure single-frame alignment quality. TNCC measures temporal coherence of consecutive full merged frames. TNCG measures deviation from $\mathrm{TNCC}_{\mathrm{ref}}=0.963$. Best in bold, second best underlined.
Method NCC $\uparrow$ SSIM $\uparrow$ TNCC $\uparrow$ TNCG $\downarrow$
Odd-only ($\mathrm{TNCC}_{\mathrm{ref}}$) 0.963
Traditional Methods
SIFT [8] 0.752 0.745 0.891 0.0726
Demons [17] 0.347 0.634 0.938 0.0259
Optical Flow [7] 0.089 0.516 0.923 0.0404
SyN [1] 0.505 0.691 0.923 0.0401
Deep Learning Methods
VoxelMorph [2] 0.741 0.594 0.893 0.0702
TransMorph [3] 0.693 0.653 0.910 0.0538
SAS-Net [10] 0.994 0.972 0.967 0.0038
PCReg-Net (Ours) 0.998 0.991 0.964 0.0001
table 2 presents the temporal evaluation results. Traditional methods that introduce frame-dependent artifacts degrade temporal coherence: SIFT (TNCG =0.0726=0.0726), Optical Flow (TNCG =0.0404=0.0404), and SyN (TNCG =0.0401=0.0401) all exhibit substantial deviation from $\mathrm{TNCC}_{\mathrm{ref}}$. Deformation-based deep learning methods similarly produce temporal artifacts: VoxelMorph (TNCG =0.0702=0.0702) and TransMorph (TNCG =0.0538=0.0538), as their learned deformation fields amplify inter-frame variability. SAS-Net achieves much better temporal consistency (TNCG =0.0038=0.0038) with strong intra-frame alignment (NCC =0.994=0.994). PCReg-Net achieves the best intra-frame quality (NCC =0.998=0.998, SSIM =0.991=0.991) while maintaining near-zero temporal consistency gap (TNCG =0.0001=0.0001), indicating that the registration preserves the natural temporal dynamics of the underlying tissue without introducing frame-to-frame artifacts.
3.4 Ablation Study
Systematic ablation experiments validate each architectural component. Five configurations are evaluated on OR-PAM-Reg-4K with 50 training epochs and seed 42.
Table 3: Ablation study on OR-PAM-Reg-4K (50 epochs). Each row removes one component from the full model. Best in bold.
Configuration NCC $\uparrow$ SSIM $\uparrow$ PSNR (dB) $\uparrow$
Full model 0.999 0.999 53.35
w/o Contrast module 0.928 0.902 32.73
w/o Feature injection 0.999 0.999 52.07
Single stage only 0.911 0.889 31.77
w/o Auxiliary loss ($\gamma=0$) 0.999 0.999 53.12
Figure 3: Ablation study results showing NCC and SSIM across five configurations. The full model achieves the best performance, with each component contributing to the final result.
Effect of contrast module.
Removing the contrast module causes a dramatic drop: NCC falls from 0.999 to 0.928 ($7.1\%$) and PSNR from 53.35 to 32.73 dB (20.62-20.62 dB). This confirms that explicitly comparing coarse-registered and reference features is critical for the refinement network to identify residual misalignment.
Effect of feature injection.
Disabling the feature injection layers reduces PSNR from 53.35 to 52.07 dB (1.28-1.28 dB) while NCC remains at 0.999. The multi-scale contrast features injected into the decoder provide complementary guidance that refines sub-pixel alignment, particularly benefiting PSNR.
Single-stage baseline.
Using only the Registration U-Net output $\hat{I}^{(c)}$ without refinement yields NCC = 0.911 and PSNR = 31.77 dB, a drop of 21.58-21.58 dB from the full model. This demonstrates that the coarse stage alone is insufficient for high-fidelity alignment, and the refinement stage provides substantial quality gains.
Effect of auxiliary loss.
Setting $\gamma=0$ slightly reduces PSNR from 53.35 to 53.12 dB (0.23-0.23 dB). While the effect is modest, auxiliary supervision on $\hat{I}^{(c)}$ provides a consistent improvement by encouraging the coarse stage to produce better initial alignment for refinement. This observation aligns with multi-task learning principles [12], where jointly optimizing related objectives yields shared representations that benefit downstream tasks.
@@ 4 Conclusion
We present PCReg-Net, a progressive contrast-guided registration framework for bidirectional OR-PAM alignment. By decomposing the registration into coarse alignment followed by contrast-guided refinement, PCReg-Net achieves high-fidelity registration quality (NCC = 0.983, SSIM = 0.982, PSNR = 46.96 dB) on the OR-PAM-Reg-4K benchmark, surpassing the state-of-the-art by over 14 dB in PSNR. Temporal consistency evaluation on OR-PAM-Reg-Temporal-26K using the proposed TNCC and TNCG metrics demonstrates that PCReg-Net preserves natural inter-frame dynamics with near-zero TNCG from the reference ceiling. The multi-scale contrast module and feature injection mechanism enable explicit identification and correction of residual misalignment at multiple scales. Ablation studies confirm the contribution of each component, and efficiency analysis demonstrates real-time capability. The framework provides a reliable solution for high-speed bidirectional OR-PAM imaging and is applicable to other scanning microscopy modalities with similar alignment challenges.

### v2

@@ 4.4 Applicability on Photoacoustic Microscopy
Unregistered (4K) 0.167 0.482 19.46
PCReg-Net (4K) 0.968 0.971 35.52
Unregistered (26K) 0.190 0.536 18.90 0.962 0.002
PCReg-Net (26K) 0.972 0.965 37.18 0.964 0.002
Odd-only ref 0.963
PCReg-Net achieves strong intra-frame registration on both OR-PAM benchmarks, with NCC of 0.968 on 4K and 0.972 on 26K. On the temporal 26K dataset, the merged frames achieve TNCC = 0.964, matching the odd-only physical reference ceiling (0.963) with minimal temporal degradation (TNCG = 0.002). These results confirm that the progressive contrast-guided refinement generalizes effectively to photoacoustic microscopy, where bidirectional scanning introduces systematic domain shift between forward and backward acquisitions.
@@ 5 Conclusion
We present PCReg-Net, a progressive contrast-guided registration framework for cross-domain image alignment. By decomposing registration into coarse alignment followed by contrast-guided refinement, PCReg-Net achieves high-fidelity registration across diverse imaging domains. The multi-scale contrast module and feature injection mechanism enable explicit identification and correction of residual misalignment at multiple scales. Comprehensive evaluation on FIRE-Reg-256 demonstrates consistent improvements over both traditional and deep learning baselines, while experiments on two OR-PAM photoacoustic microscopy benchmarks confirm cross-domain generalization with strong registration quality. Ablation studies confirm the contribution of each architectural component, with the contrast module providing the most significant gain. With only 2.56M parameters and 141 FPS inference speed, PCReg-Net offers an efficient and generalizable solution for cross-domain image registration.